(c)
$P(A \cup B) = P(A) + P(B) - P(A \cap B)$
$\therefore 0.6 = P(A) + P(B) - 0.2$
$\Rightarrow P(A) + P(B) = 0.8$
$\Rightarrow P\left( \overline{A} \right) + P\left( \overline{B} \right) = 12\ \ \lbrack\because P(A) = 1 - P\left( \overline{A} \right)\rbrack$